OpenStudy (anonymous):

if an aroma of vanillin fills a room, having the dimensions of 12ftx10.5ftx10.0ft. if the density is 2.4x10^-11g/L how many kilograms of vanillin are in the room?

sam (.sam.):

Density formula: \[\rho=\frac{m}{V}\] \[m=\rho V\] ------------------------------------------------------- Volume= 12x10.5x10.0=1260ft^3= 35679 liters \[m=\rho V\] \[m=( 2.4x10^{-11}g/L )(35679L) \] \[m=8.56\times 10^{-7}g=8.56 \times 10^{-10}kg\]

OpenStudy (anonymous):

hey thanks- can you explain how you get from cubic feet to liters?

sam (.sam.):

You either remembering it or from a table of conversions, 1 Cubic Feet = 28.32 Liters You can arrange it so that the cubic feet is at the bottom when you multiply, to cancel the cubic feet. You want liters here, so: \[1280~ ft^3 \times \frac{28.32~liters}{1~Cubic~feet}=35679~liters\]

OpenStudy (anonymous):

got it. thank you so much!